OpenStudy (anonymous):

Determine the slope of (a) every line that is parallel to the given line and (b) every line 2y-5x+4=0 Already helped finding the slope that is parallel as 5/2, but i still cannot figure out the line that is perpendicular to the line...unless it's as simple as -5/2

OpenStudy (anonymous):

2y - 5x + 4 = 0 2y = 5x - 4 y = 5/2x - 4 you are correct parallel is 5/2 perpendicular is the negative reciprocal so perpendicular is -2/5